OpenStudy (jacky90):

HELP please will medal There are 810 dishes that need to be rinsed. Miguel can rinse them in 135 minutes by himself. It will take his friend Ravi 90 minutes to rinse these dishes. How long will it take them if they rinse these 810 dishes together?

OpenStudy (anonymous):

John`s rate of work = 810 / 135 dishes / min Bob`s rate of work = 810 / 90 dishes / min Combined rate = 810/135 + 810/90 dishes / min Combined rate = 6 + 9 = 15 dishes / min Time = 810 / 15 min = 54 min